# Doing Experiments and Revising Rules with Natural Language and Probabilistic Reasoning

### v1

@@ Appendix A Appendix
Batch Inference (with and without Rule Refinement) with Hard Rules
{spverbatim}
Given the following structures described with att_summary of blocks in the structures: text_c Please list num possible rules about the attributes in a structure that differentiate the good structures from the bad structures. Keep in mind that 1. All bad structures must violate the rules. 2. Orders of blocks in a structure do NOT matter. 3. The rules are short, concise, single sentences. Please number them from 1-num and do not say anything else
Online Inference with Hard Rules
{spverbatim}
Please list num possible rules about the att att_choices.
Example 1: Structure: blue, blue Simple rules (Orders do NOT matter): 1. There is a blue block 2. There is no green block 3. All blocks are blue
Task 1: x Simple rules (Orders do NOT matter):
Batch Inferece with Fuzzy Rules
{spverbatim}
Given the following structures described with att_summary of blocks in the structures: text_c Please list num possible rules about the attributes in a structure that differentiate the good structures from the bad structures. Keep in mind that 1. All bad structures must violate the rules. 2. Orders of blocks in a structure do NOT matter. 3. Do NOT propose ”negative rules” such as ”there is no green block”. 4. The rules are short, concise, single sentences. Please number them from 1-num and do not say anything else
Online Inference with Fuzzy Rules
{spverbatim}
Please list num possible rules about the att.
Example 1: Structure: blue, blue Simple rules (Orders do NOT matter): 1. There is a blue block 2. All blocks are blue Do NOT propose ”negative rules” such as ”there is no green block”. Do NOT propose rules with quantifier such as ”there are two blue blocks”
Task 1: x Simple rules (Orders do NOT matter):
@@
{spverbatim}
A structure has one or more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch).
Example of rule modifications: Quantifier change: ’There must be a green block’ -¿ ’There are two green blocks’ Additional attribute: ’There must be a green block’ -¿ ’There must be a green block that is upright’ Attribute change: ’There must be a green block’ -¿ ’There must be a blue block’ These modifications are ”local”: only one attribute/quantifier is changed or added for each modification.
Please modify the rule ’h’. Generate num rules for each type of modification (Quantifier change, Additional attribute, Attribute change) so that the following structure is text_y a good structure: x Note that the number of the blocks do not matter.
Make the format a numbered list (1., 2., …, 15.) Remember that the new rules should be a ”local” modification from the rule ’h’. Do not use attribute values that are not mentioned earlier. Do not say anything other than the modified rules.
@@
{spverbatim}
A structure has one or more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch).
Consider the following rule: ’h’
Given a structure, the output is yes if it follows the rule (or is a good structure) and no if it does not (or is a bad structure)
The given rule gives incorrect output for the following structures:
feedback
Based on the given rule, generate num new refined rules that fix the outputs for all mentioned structures. The new rules may involve any of the mentioned attributes (color, size, orientation, grounded, touching). Please number them from 1-num and do not say anything else
@@
Initial prompt
{spverbatim}
You are playing an inductive game with me. I’ll be the moderator, and your task is to figure out the secret rule that I know by coming up with a structure of blocks to ask me whether it conforms with the secret rule or not.
The structure has one of more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch).
To give you a start, I’ll describe one structure that follows the rule:
text_c
Give a very short summary on what you currently think the secret rule is.

### v3

@@ Introduction
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ Appendix A Appendix
Batch Inference (with and without Rule Refinement) with Hard Rules {spverbatim} Given the following structures described with att_summary of blocks in the structures: text_c Please list num possible rules about the attributes in a structure that differentiate the good structures from the bad structures. Keep in mind that 1. All bad structures must violate the rules. 2. Orders of blocks in a structure do NOT matter. 3. The rules are short, concise, single sentences. Please number them from 1-num and do not say anything else
Online Inference with Hard Rules {spverbatim} Please list num possible rules about the att att_choices. Example 1: Structure: blue, blue Simple rules (Orders do NOT matter): 1. There is a blue block 2. There is no green block 3. All blocks are blue Task 1: x Simple rules (Orders do NOT matter):
Batch Inferece with Fuzzy Rules {spverbatim} Given the following structures described with att_summary of blocks in the structures: text_c Please list num possible rules about the attributes in a structure that differentiate the good structures from the bad structures. Keep in mind that 1. All bad structures must violate the rules. 2. Orders of blocks in a structure do NOT matter. 3. Do NOT propose ”negative rules” such as ”there is no green block”. 4. The rules are short, concise, single sentences. Please number them from 1-num and do not say anything else
Online Inference with Fuzzy Rules {spverbatim} Please list num possible rules about the att. Example 1: Structure: blue, blue Simple rules (Orders do NOT matter): 1. There is a blue block 2. All blocks are blue Do NOT propose ”negative rules” such as ”there is no green block”. Do NOT propose rules with quantifier such as ”there are two blue blocks” Task 1: x Simple rules (Orders do NOT matter):
@@
{spverbatim} A structure has one or more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch). Example of rule modifications: Quantifier change: ’There must be a green block’ -¿ ’There are two green blocks’ Additional attribute: ’There must be a green block’ -¿ ’There must be a green block that is upright’ Attribute change: ’There must be a green block’ -¿ ’There must be a blue block’ These modifications are ”local”: only one attribute/quantifier is changed or added for each modification. Please modify the rule ’h’. Generate num rules for each type of modification (Quantifier change, Additional attribute, Attribute change) so that the following structure is text_y a good structure: x Note that the number of the blocks do not matter. Make the format a numbered list (1., 2., …, 15.) Remember that the new rules should be a ”local” modification from the rule ’h’. Do not use attribute values that are not mentioned earlier. Do not say anything other than the modified rules.
@@
{spverbatim} A structure has one or more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch). Consider the following rule: ’h’ Given a structure, the output is yes if it follows the rule (or is a good structure) and no if it does not (or is a bad structure) The given rule gives incorrect output for the following structures: feedback Based on the given rule, generate num new refined rules that fix the outputs for all mentioned structures. The new rules may involve any of the mentioned attributes (color, size, orientation, grounded, touching). Please number them from 1-num and do not say anything else
@@
Initial prompt {spverbatim} You are playing an inductive game with me. I’ll be the moderator, and your task is to figure out the secret rule that I know by coming up with a structure of blocks to ask me whether it conforms with the secret rule or not. The structure has one of more blocks. Each block should contain the following attributes: color (blue/red/green) size (small/medium/large) orientation (upright/left/right/strange) groundedness (grounded/ungrounded), touching (which other blocks they do and do not touch). To give you a start, I’ll describe one structure that follows the rule: text_c Give a very short summary on what you currently think the secret rule is.